# The sup-norm problem for newforms of large level on $\operatorname{PGL}(n)$

### v1

@@ 5. Reduction of the domain
The value of $\phi(z)$ is independent of which element in the orbit $\Gamma_{0}(N)\cdot z$ we choose instead of $z$. Similarly, the number and shape modulo $N$ of the matrices we are considering in the amplified pre-trace formula in Proposition 3.3 is invariant under shifting by elements of $\Gamma_{0}(N)$, which would merely amount to conjugating $H(z,m,N)$.
Consider now the action of the Fricke involution $W_{N}(\phi)(z)=\phi(z^{\prime})$. If $Y\subset\Gamma_{0}(N)$ is a subset, we denote by $Y^{\prime}$ the image of $Y$ under the map $z\mapsto z^{\prime}$. It is clear that we obtain a bound for a Hecke-Maaß form $\phi$ on $Y\cup Y^{\prime}$ if we have a bound for both $\phi$ and $W_{N}(\phi)$ on the subset $Y$.
@@
$\Gamma_{0}(N)z\cup\Gamma_{0}(N)z^{\prime}$
@@
In this section we propose a system for making this selection of representative. In other words, we construct an approximate fundamental domain for the action of $\Gamma_{0}(N)$ and the Fricke involution, at least in the bulk. It can be seen as a reduction theory with level structure, for which we often use the shorter term Fricke reduction.
@@ 5.1. Two lattices
Throughout the following sections we assume that $N$ is a prime.
@@
$z^{\prime}:=A_{N}z^{-T}=N^{-1/n}\operatorname{diag}(1,\ldots,1,N)z^{-T}.$
@@
$A(z)$ =$=\{\left\lVert e_{n}\right\rVert_{\gamma z}\mid\gamma\in\Gamma_{0}(N)\},$
$B(z)$ =$=\{\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{\gamma z}\mid\gamma\in\Gamma_{0}(N)\}.$
In the following paragaphs we show how the union of $A(z),B(z),A(z^{\prime}),B(z^{\prime})$ provides the lengths of all primitive vectors in $L_{z}$, $L_{z^{\prime}}$, and their duals.
@@
$\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{\gamma z^{\prime}}=\left\lVert e_{1}\right\rVert_{\gamma^{-T}A_{N}^{-1}z}=N^{1/n}\left\lVert(a_{1},\ldots,a_{n})\right\rVert_{z},$ (5.1)
@@
$\gcd(\gcd(a_{1},\ldots,a_{n-1}),N)=1.$
Since $N$ is prime, the greatest common divisor of $\gcd(a_{1},\ldots,a_{n-1})$ and $N$ can only be $1$ or $N$, and thus, considering the paragraph above we have exhausted all primitive vectors in $L_{z}$.
@@
where $h\in\operatorname{SL}_{n-1}(\mathbb{Z})$, leaves the first row invariant. We shall inductively apply such row operations on $g$ to make its last column be of the form $(c_{1},\ldots,c_{n})$, where $c_{1}=Na_{n}$ and $N$ divides $c_{1},\ldots,c_{n-1}$. Indeed, if $N\mid c_{i}$ for any $i\in\{2,\ldots,n\}$, then we can permute rows to assume that $N\mid c_{2}$. Otherwise we can assume that $\gcd(N,c_{3})=1$. Let $\bar{c}_{3}$ be any representative of the inverse of $c_{3}$ modulo $N$. Bézout’s lemma provides a matrix $h^{\prime}\in\operatorname{SL}_{2}(\mathbb{Z})$ with top row $(N,\bar{c}_{3})$. Using $h$ of the form
@@
$\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{\gamma z}=\left\lVert e_{1}\right\rVert_{\gamma^{-T}z^{-T}}=\left\lVert(a_{1},\ldots a_{n-1},Na_{n})\right\rVert_{z^{-T}},$ (5.2)
@@
$\left\lVert e_{n}\right\rVert_{\gamma A_{N}z^{-T}}=N^{1-1/n}\left\lVert(a_{1},\ldots,a_{n})\right\rVert_{z^{-T}},$ (5.3)
@@ 5.2. Fricke reduction
Let $z\in\mathbb{H}$. For the study of the sup-norm and our counting problem, we are allowed to choose any conjugate of $z$ in the orbit $\Gamma_{0}(N)\cdot z$ and also switch between $z$ and $z^{\prime}$, as explained at the beginning of Section 5. Now it is clear by construction that every $z$ is contained in some $\mathcal{L}(X,Y)$. We then make the choice of conjugate to obtain a well-positioned $z$, where we have control over its successive minima and Iwasawa coordinates, based on which set $\mathcal{L}(X,Y)$ contains $z$.
@@ 5.2.1. Case I
In particular, $l\gg\lambda_{1}^{2}$. This implies that
@@
[…]
@@ 5.2.2. Case II
$z\in[\mathcal{L}(B,B)\cap\mathcal{L}^{\prime}(B,A)]\cup[\mathcal{L}(B,A)\cap\mathcal{L}^{\prime}(B,B)]\cup[\mathcal{L}(B,B)\cap\mathcal{L}^{\prime}(B,B)].$
Applying the Fricke involution if needed, we can assume that $z$ lies in $\mathcal{L}^{\prime}(B,B)$ and in $\mathcal{L}(B,\ast)$. Then the minimal length in $L^{\ast}_{z^{\prime}}$ is given by $b(z^{\prime})$ and the minimal length in $L_{z}$ is $N^{-1/n}b(z^{\prime})$.
By Minkowski’s theorem, more precisely equation (2.2), applied to $L_{z^{\prime}}^{\ast}$, we find that $b(z^{\prime})\ll_{n}1$. This now implies that the minimal length in $L_{z}$ is $N^{-1/n}b(z^{\prime})\ll N^{-1/n}$.
5.2.3. Case III
Let
$z\in\mathcal{L}(B,A)\cap\mathcal{L}^{\prime}(B,A).$
Applying the Fricke involution if required, we may assume that $\alpha(z^{\prime})\leq\alpha(z)$. Shifting $z=n(x)a(y)$ by a suitable $\gamma\in\Gamma_{0}(N)$ as in Case I, i.e. Section 5.2.1, we also assume that $\alpha(z)=\left\lVert e_{n}\right\rVert_{z}=d$ and that $y_{i}\gg 1$ for $i=2,\ldots,{n-1}$.
Note now that the minimal lenght in $L_{z^{\prime}}^{\ast}$ is $N^{-1+1/n}\alpha(z)$. Note also that $L_{z^{\prime}}^{\ast}$ is the lattice corresponding to
@@ 5.2.3. Case III
Using Lemma 2.2, we deduce that the minimal lenght $l$ in $\bigwedge^{2}L_{z^{\prime}}^{\ast}$ satisfies $l\gg\mu_{1}^{2}$, where $\mu_{1}$ is the first successive minimum of $L_{z^{\prime}}^{\ast}$. Putting everything together we arrive at
@@
which implies that $y_{1}\gg_{n}N^{-1}$.
@@ 5.2.4. Fricke reduction of points that reduce to a compactum
We summarise the cases described above in the context of points $z$ that reduce to a fixed compact set $\Omega\subset\mathbb{H}$.
@@
Let $z\in\mathbb{H}$ and let $\Omega\subset\mathbb{H}$ be a compact set. For $N\gg_{\Omega}1$ prime, large enough, there is
$w\in\{\gamma z\mid\gamma\in\Gamma_{0}(N)\}\cup\{\gamma z^{\prime}\mid\gamma\in\Gamma_{0}(N)\},$
where $z^{\prime}=A_{N}z^{-T}$, with Iwasawa coordinates $w=n(x)a(y)$ which, if $z$ reduces to $\Omega$, satisfy either
$y_{i}\asymp_{\Omega}1$
for all $i=1,\ldots,n-1$, in which case
$w\in\bigcup_{X\in\{A,B\}}\mathcal{L}(A,X)\cup\mathcal{L}^{\prime}(A,X)$
or
[…]
for $i=2,\ldots,n-1$, in which case
@@
Proof.
By Lemma 2.7, we eliminate Case II, since there the minimal length in $L_{z}$ is $\ll N^{-1/n}$. In Case I we find $w$ as in the statement such that $y_{i}\gg_{n}1$ for all $i=1,\ldots,n-1$. Thus $w$ lies in a Siegel set and Lemma 2.5 together with Lemma 2.7 implies that $y_{i}\asymp_{n,\Omega}1$ for all $i$. In Case III we find $w$ such that $\alpha(w^{\prime})\leq\alpha(w)$, $y_{1}\gg_{n}N^{-1}$, and $y_{i}\gg 1$ for $i=2,\ldots,n-1$. Since $w\in\mathcal{L}(B,A)$, the minimal length in $L^{\ast}_{w}$ is $N^{-1+1/n}\alpha(w^{\prime})$ and by Lemma 2.7 we deduce that $\alpha(w^{\prime})\gg_{\Omega}N^{1-1/n}$. Since $\alpha(w^{\prime})\leq\alpha(w)$, we also have that $\alpha(w)\gg N^{1-1/n}$. Now $\alpha(w)=\left\lVert e_{n}\right\rVert_{z}=d$. Writing out the definition of $d$, we see that
@@
Combining this with the bounds above for the $y$-coordinates, we deduce that $y_{1}\asymp N^{-1}$ and $y_{i}\asymp 1$ for $i=2,\ldots,n-1$, where the implicit constants depend on $\Omega$. ∎

### v2

@@ 5. Reduction of the domain
For some $w\in\mathbb{H}^{n}$, the value of $\phi(w)$ is independent of which element in the orbit $\Gamma_{0}(N)\cdot w$ we choose instead of $w$. Similarly, the number and shape modulo $N$ of the matrices we are considering in the amplified pre-trace formula in Proposition 3.3 is invariant under shifting by elements of $\Gamma_{0}(N)$, which would merely amount to conjugating $H(w,m,N)$.
Consider now the action of the Fricke involution $W_{N}(\phi)(w)=\phi(w^{\prime})$. If $Y\subset\Gamma_{0}(N)$ is a subset, we denote by $Y^{\prime}$ the image of $Y$ under the map $w\mapsto w^{\prime}$. It is clear that we obtain a bound for a Hecke-Maaß form $\phi$ on $Y\cup Y^{\prime}$ if we have a bound for both $\phi$ and $W_{N}(\phi)$ on the subset $Y$.
@@
$\Gamma_{0}(N)w\cup\Gamma_{0}(N)w^{\prime}$
@@
In this section we propose a system for making this selection of representative. In other words, we construct an approximate fundamental domain for the action of $\Gamma_{0}(N)$ and the Fricke involution, at least on $\Omega_{N}^{\natural}$. It can be seen as a reduction theory with level structure, for which we often use the shorter term Fricke reduction.
5.1. A reinterpretation of the Harcos–Templier reduction
To motivate our approach, we first revisit the case of $\operatorname{PGL}(2)$, assuming that $N$ is prime. Here, the elegant and simple idea of Harcos–Templier [21] is to choose a representative
$z\in\Gamma_{0}(N)w\cup\Gamma_{0}(n)w^{\prime}$
such that $\Im z$ is maximal when viewing $z$ in the upper-half plane. Using the Iwasawa coordinates $z=n(x)a(y)$, where $a(y)=d\cdot\operatorname{diag}(y,1)$, this amounts to minimising $\left\lVert e_{2}\right\rVert_{z}$. What this achieves is that $y\gg 1/N$ and, in the upper-half plane, $|cz+d|\geq 1/\sqrt{N}$ for all $(c,d)\in\mathbb{Z}^{2}$ (see Lemma 1 in [21]). After normalisation, the latter condition is equivalent to saying the shortest length in $L_{z}$ is bounded from below by $d/\sqrt{N}$. Note that $d=1/\sqrt{y}$.
This last bound shows that $y$ controls the balancedness of $L_{z}$, which plays a crucial role in the counting problem. For $y$ close to $1/N$, the lattice $L_{z}$ is balanced and this is tractable, yet for larger $y$ they resort to applying a Fourier bound.
We now look closer at this reduction process and compare the two lattices $L_{w}$ and $L_{w^{\prime}}$. Define
$\alpha(w)=\min\{\left\lVert e_{2}\right\rVert_{\gamma z}\mid\gamma\in\Gamma_{0}(N)\}$
and similarly for $\alpha(w^{\prime})$. Now $e_{2}\cdot\Gamma_{0}(N)$ does not exhaust all primitive vectors in $\mathbb{Z}^{2}$ and thus $\alpha(w)$ does not necessarily give the first successive minimum of $L_{w}$. However, since $N$ is prime, we only lack primitive vectors of the form $(a,b)$, where $\gcd(a,N)=1$. In this case $(Nb,a)\in\mathbb{Z}^{2}$ is another primitive vector. We may thus find $\gamma\in\Gamma_{0}(N)$ with last row equal to $(Nb,a)$.
Consider, for simplicity, the old definition of the Fricke involution
$w^{\prime}=N^{-1/2}\begin{pmatrix}0&-1\\ N&0\end{pmatrix}w.$
A quick computation shows that
$\left\lVert e_{2}\right\rVert_{\gamma w^{\prime}}=\left\lVert(Nb,a)\right\rVert_{w^{\prime}}=\sqrt{N}\left\lVert(a,b)\right\rVert_{w}.$
Putting everything together, the shortest length in $L_{w}$ is either $\alpha(w)$ or $\alpha(w^{\prime})/\sqrt{N}$. By symmetry, the shortest length in $L_{w^{\prime}}$ is either $\alpha(w^{\prime})$ or $\alpha(w)/\sqrt{N}$.
To choose a representative $z$, what Harcos–Templier do is to potentially apply the Fricke involution to achieve that $\alpha(z)\leq\alpha(z^{\prime})$ and then to shift $z$ by $\Gamma_{0}(N)$ to achieve $\alpha(z)=d$, where $d=1/\sqrt{\Im z}$. We now look at what information this choice implies about the lattices above, and we distinguish a few possibilities.
If $\alpha(z)=\left\lVert e_{2}\right\rVert_{z}=d$ is the shortest length in $L_{z}$, then we have achieved the first step in Gauß–Lagrange reduction, and we may continue to obtain $y\geq\sqrt{3}/2$. For this conclusion we do not use that $\alpha(z)\leq\alpha(z^{\prime})$. In fact, since this lower bound on $y$ is so strong, it seems useful to prioritise using the Fricke involution to arrive at this situation. Thus, in the case that $\alpha(z^{\prime})$ is the shortest length in $L_{z^{\prime}}$, we could apply the Fricke involution and obtain the same conclusion.
We are left with the case where the shortest length in $L_{z}$ is $\alpha(z^{\prime})/\sqrt{N}$ and the shortest in $L_{z^{\prime}}$ is $\alpha(z)/\sqrt{N}$. We assume the Harcos–Templier conditions and investigate what happens if we assume that our original point reduces to a compactum $\Omega$. Interpreting this in terms of balancedness of lattices but not knowing whether the Fricke involution was applied or not, we must distinguish two cases.
First, if we applied the involution and $z^{\prime}$ reduces to $\Omega$, meaning that $L_{z^{\prime}}$ is balanced, then its shortest length $\alpha(z)/\sqrt{N}$ is approximately $1$, implying that $d=1/\sqrt{y}\asymp_{\Omega}\sqrt{N}$. Thus, we have the case $y=1/N$, where the counting results work best. As noted before, this also implies that $L_{z}$ is balanced. Indeed, the shortest length in $L_{z}$ is $\alpha(z^{\prime})/\sqrt{N}\geq\alpha(z)/\sqrt{N}\asymp 1$, by assumption.
Otherwise, if $L_{z}$ is the balanced lattice and we cannot assume anything about $L_{z^{\prime}}$, notice that the crucial parameter $y$ still controls the balancedness of $L_{z^{\prime}}$. Thus, either $L_{z^{\prime}}$ is rather balanced, or $y$ is much larger than $1/N$. This dichotomy can be successfully used in solving the sup-norm problem, as long as the Fourier bound is effective for $y\gg N^{-1+\delta}$ for small enough $\delta$.
The case of $y$ much larger than $1/N$ seems possible, since no contradictions can be found with our assumptions. In fact, $L_{z}$ being balanced implies the trivial bound $\alpha(z)\gg 1$, translating only to $y\ll 1$. For instance, this shows that the shortest length in $L_{z^{\prime}}$ is bounded from below by roughly $1/\sqrt{N}$.
The considerations above show that the preimage of a compact set $\Omega$ under the natural projection map $X_{2}(N)\to X_{2}(1)$ can be covered by $Y\cup Y^{\prime}$, where
$Y\subset\{z=x+iy\mid x\in[-1/2,1/2],y\gg 1/N\}.$
Harcos–Templier show the same for the full preimage of $X_{2}(1)$, yet the inclusion above seems sharp even for a compact set.
@@ 5.2. Two lattices
We now extend the ideas above to higher rank, which is significantly more complicated. Throughout the following sections we assume that $N$ is a prime.
@@
$z^{\prime}:=A_{N}w^{-T}=N^{-1/n}\operatorname{diag}(1,\ldots,1,N)z^{-T}.$
@@
$A(z)$ $=\{\left\lVert e_{n}\right\rVert_{\gamma z}\mid\gamma\in\Gamma_{0}(N)\},$
$B(z)$ $=\{\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{\gamma z}\mid\gamma\in\Gamma_{0}(N)\}.$
In the following paragraphs we show how the union of $A(z),B(z),A(z^{\prime}),B(z^{\prime})$ provides the lengths of all primitive vectors in $L_{z}$, $L_{z^{\prime}}$, and their duals.
@@
(5.1) $\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{\gamma z^{\prime}}=\left\lVert e_{1}\right\rVert_{\gamma^{-T}A_{N}^{-1}z}=N^{1/n}\left\lVert(a_{1},\ldots,a_{n})\right\rVert_{z},$
@@
$\gcd(a_{1},\ldots,a_{n-1},N)=1.$
Since $N$ is prime, the greatest common divisor of $a_{1},\ldots,a_{n-1}$ and $N$ can only be $1$ or $N$. Thus, considering the paragraph above, we have exhausted all primitive vectors in $L_{z}$.
@@
where $h\in\operatorname{SL}_{n-1}(\mathbb{Z})$, leaves the first row invariant. We shall inductively apply such row operations on $g$ to make its last column be of the form $(c_{1},\ldots,c_{n})$, where $c_{1}=Na_{n}$ and $N$ divides $c_{1},\ldots,c_{n-1}$.
Indeed, if $N\mid c_{i}$ for any $i\in\{2,\ldots,n\}$, then we can permute rows to assume that $N\mid c_{2}$. Otherwise, we can assume that $\gcd(N,c_{3})=1$. Let $\bar{c}_{3}$ be any representative of the inverse of $c_{3}$ modulo $N$. Bézout’s lemma provides a matrix $h^{\prime}\in\operatorname{SL}_{2}(\mathbb{Z})$ with top row $(N,\bar{c}_{3})$. Using $h$ of the form
@@
(5.2) $\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{\gamma z}=\left\lVert e_{1}\right\rVert_{\gamma^{-T}z^{-T}}=\left\lVert(a_{1},\ldots a_{n-1},Na_{n})\right\rVert_{z^{-T}},$
@@
(5.3) $\left\lVert e_{n}\right\rVert_{\gamma A_{N}z^{-T}}=N^{1-1/n}\left\lVert(a_{1},\ldots,a_{n})\right\rVert_{z^{-T}},$
@@ 5.3. Fricke reduction
Let $w\in\mathbb{H}$. For the study of the sup-norm and our counting problem, we are allowed to choose any conjugate of $z$ in the orbit $\Gamma_{0}(N)\cdot w$ and also switch between $z$ and $z^{\prime}$, as explained at the beginning of Section 5. Now, it is clear by construction that every $w$ is contained in some $\mathcal{L}(X,Y)$. We then make the choice of conjugate to obtain a well-positioned $z$, where we have control over its successive minima and Iwasawa coordinates, based on which set $\mathcal{L}(X,Y)$ contains $z$. The guiding principle is that $y$-coordinates of $z$ should be bounded from below and as large as possible, similar to classical reduction theory. If this is not possible, we try to obtain some inbalancedness information on the relevant lattices.
We first begin with some preliminary general observations and then construct a domain in the following proposition, under a balancedness assumption.
@@ 5.3.1. Case I
In particular, $l\gg\lambda_{1}^{2}$. This implies that (recall the calculation in Remark 2.4)
@@
We deduce that $y_{1}\gg_{n}1$. [^6] This can be viewed as a soft version of Hermite reduction, that is, reduction to a Siegel set. Indeed, here we also take the last row to be the shortest vector and then use induction, as in the classical proof of reduction.
@@ 5.3.2. Case II
$z\in[\mathcal{L}(B,A)\cap\mathcal{L}^{\prime}(B,A)]\cup[\mathcal{L}(B,B)\cap\mathcal{L}^{\prime}(B,A)].$
Shifting $z=n(x)a(y)$ by a suitable $\gamma\in\Gamma_{0}(N)$ as in Case I, we assume that $\alpha(z)=\left\lVert e_{n}\right\rVert_{z}=d$ and that $y_{i}\gg 1$ for $i=2,\ldots,{n-1}$.
Note now that the minimal length in $L_{z^{\prime}}^{\ast}$ is $N^{-1+1/n}\alpha(z)$. Note also that $L_{z^{\prime}}^{\ast}$ is the lattice corresponding to
@@
Using Lemma 2.2, we deduce that the minimal length $l$ in $\bigwedge^{2}L_{z^{\prime}}^{\ast}$ satisfies $l\gg\mu_{1}^{2}$, where $\mu_{1}$ is the first successive minimum of $L_{z^{\prime}}^{\ast}$. Putting everything together we arrive at
@@
which implies that $y_{1}\gg_{n}N^{-1}$. Notice that the Fricke involution has not been used.
5.3.3. Case III
Let
$z\in[\mathcal{L}(B,A)\cap\mathcal{L}^{\prime}(B,B)]\cup[\mathcal{L}(B,B)\cap\mathcal{L}^{\prime}(B,B)].$
Then the minimal length in $L^{\ast}_{z^{\prime}}$ is given by $b(z^{\prime})$ and the minimal length in $L_{z}$ is $N^{-1/n}b(z^{\prime})$. By Minkowski’s theorem, more precisely equation (2.2), applied to $L_{z^{\prime}}^{\ast}$, we find that $b(z^{\prime})\ll_{n}1$. This now implies that the minimal length in $L_{z}$ is $N^{-1/n}b(z^{\prime})\ll N^{-1/n}$. Notice that we again do not need the Fricke involution for this conclusion.
@@ 5.3.4. Fricke reduction of points that reduce to a compactum
We summarise the cases described above in the context of points $w$ that reduce to a fixed compact set $\Omega\subset\mathbb{H}$ and we produce a well-positioned conjugate $z$.
@@
Let $w\in\mathbb{H}^{n}$, let $\Omega\subset\mathbb{H}^{n}$ be a compact set and assume that $w$ reduces to $\Omega$. For $N\gg_{\Omega}1$ prime, large enough, only Case I and Case II are valid. There is
$z\in\{\gamma w\mid\gamma\in\Gamma_{0}(N)\}\cup\{\gamma w^{\prime}\mid\gamma\in\Gamma_{0}(N)\},$
where $w^{\prime}=A_{N}w^{-T}$, with Iwasawa coordinates $z=n(x)a(y)$,
$a(y)=d\cdot\begin{pmatrix}y_{1}\cdots y_{n-1}&&&\\ &\ddots&&\\ &&y_{1}&\\ &&&1\end{pmatrix}$,
which satisfies:
  1. (1)
    if $w$ is in Case I, then
    $y_{i}\gg 1\text{ for all }i=1,\ldots,n-1;$
  2. (2)
    if $w$ is in Case II, then we can arrange that $d=\alpha(z)$, we have the lower bounds
    $y_{1}\gg\frac{1}{N}\quad\text{and}\quad y_{i}\gg 1\text{ for }i=2,\ldots,n-1,$
    and either
    1. (a)
      $w$ belongs to a set of volume $\ll N^{(n-1)(1-1/2n^{4})}$ and we have
      $y_{1}\gg N^{-1+1/2n^{4}}\text{ or }y_{i}\gg N^{1/2n^{4}}\text{ for some }2\leq i\leq n-1;$
    2. (b)
      or $z,w\in\mathcal{L}(B,A)\cap\mathcal{L}^{\prime}(B,A)$, the lattice $L_{z}$ is $\Omega$-balanced, we have
      $\lambda_{1}(L_{z^{\prime}})\gg N^{-1/2n}$
      and we have the upper bounds
      $dy_{1}\cdots y_{k}\ll N^{-1/2n}\text{ for any }1\leq k\leq n-1.$
All implied constants may depend on $\Omega$ and $n$.
Proof.
In Case I, we have already shown how to find $z$ as in the statement such that $y_{i}\gg_{n}1$ for all $i=1,\ldots,n-1$.
Case II is more delicate and we first consider the possibility that
$w\in\mathcal{L}(B,B)\cap\mathcal{L}^{\prime}(B,A).$
Here, as explained above, we may find $z=\gamma w$ with $\gamma\in\Gamma_{0}(N)$ with Iwasawa coordinates satisfying that $d=\alpha(z)$, $y_{1}\gg N^{-1}$, and $y_{i}\gg 1$ for $i\geq 2$. Observe now that the shortest vector in $L_{z}^{\ast}$ is given by $\beta(z)$. Since we did not apply the Fricke involution, $z$ reduces to $\Omega$. By Lemma 2.8, $L_{z}^{\ast}$ is a balanced lattice and therefore $\beta(z)\asymp_{\Omega}1$.
On the other hand, the shortest vector in $L_{z^{\prime}}$ is given by $N^{-1/n}\beta(z)$ and this is now of size $N^{-1/n}$. By Minkowski’s theorem, we have $\lambda_{1}(L_{z^{\prime}})\lambda_{n}(L_{z^{\prime}})^{n-1}\gg 1$ and therefore $\lambda_{n}(L_{z^{\prime}})\gg N^{1/n(n-1)}$. By a standard transference theorem, putting together Lemma 2.1 and Lemma 2.2, we find that
$\lambda_{1}(L_{z^{\prime}}^{\ast})\ll\lambda_{n}(L_{z^{\prime}})^{-1}\ll N^{-1/n(n-1)}.$
Since the shortest vector in $L_{z^{\prime}}^{\ast}$ is given by $N^{-1+1/n}\alpha(z)$, our assumptions imply that $d\ll N^{1-1/n-\eta_{n}}$, where $\eta_{n}=1/n(n-1)$.
Since $d^{n}y_{1}^{n-1}\cdots y_{n-1}=1$ by assumption, we deduce from the upper bound on $d$ that $y_{1}\gg N^{-1+\eta_{n}^{\prime}}$ or $y_{i}\gg N^{\eta_{n}^{\prime}}$ for some $i\geq 2$, where we may take $\eta_{n}^{\prime}=\eta_{n}/(n-1)^{2}$. We may bound now the volume of the Siegel-like set with such lower bounds on the $y$-coordinates (and $x$-coordinates uniformly bounded, as usual). Recall (see [20, Prop. 1.5.3]) that the volume form for $\mathbb{H}_{n}$ in the $y$-coordinates is
$\prod_{k=1}^{n-1}y_{k}^{-k(n-k)-1}\,dy_{k}.$
The largest exponent appearing is n-n and therefore the bounds above show that the volume is bounded by $N^{(n-1)(1-\eta_{n}^{\prime})}$.
We now consider the crucial case of
@@
This case is invariant under the Fricke involution, and we apply it if necessary to obtain $z$ such that $\alpha(z)\leq\alpha(z^{\prime})$. Shifting by $\Gamma_{0}(N)$, we may also assume that the $y$-coordinates of $z$ satisfy $d=\alpha(z)$, $y_{1}\gg_{n}N^{-1}$, and $y_{i}\gg 1$ for $i=2,\ldots,n-1$, as we generally do with Case II.
Since we potentially applied the Fricke involution, we must now consider two possibilities. First, if we did apply the involution and it is $z^{\prime}$ that reduces to $\Omega$, then $L_{z^{\prime}}$ and $L_{z^{\prime}}^{\ast}$ are $\Omega$-balanced. Since $z\in\mathcal{L}^{\prime}(B,A)$, the minimal length in $L^{\ast}_{z^{\prime}}$ is $N^{-1+1/n}\alpha(z)$ and we deduce that $\alpha(z)=d\asymp N^{1-1/n}$. We see that
@@
Combining this with the bounds above for the $y$-coordinates, we deduce that $y_{1}\asymp N^{-1}$ and $y_{i}\asymp 1$ for $i=2,\ldots,n-1$, where the implicit constants depend on $\Omega$. Additionally, $z\in\mathcal{L}(B,A)$ implies that the shortest length in $L_{z}^{\ast}$ is
$N^{-1+1/n}\alpha(z^{\prime})\geq N^{1+1/n}\alpha(z)\gg 1.$
As such, $L_{z}^{\ast}$ is a balanced lattice and, by Lemma 2.8, the same is true about $L_{z}$.
Otherwise, we have that $z$ reduces to $\Omega$, implying that $L_{z}$ is balanced. In this case, we cannot deduce that $L_{z^{\prime}}$ is balanced, but any imbalancedness translates into lower bounds for $y$-coordinates, as above.
Indeed, by assumption, the shortest vector in $L_{z^{\prime}}$ has length $N^{-1/n}\beta(z)$. If $N^{-1/n}\beta(z)\leq N^{-1/2n}$, then the same arguments as above show that
$d\ll N^{1-1/n-1/2n(n-1)},\quad y_{1}\gg N^{-1+1/2n(n-1)^{3}},\quad y_{i}\gg N^{1/2n(n-1)^{3}}\text{ for all }i\geq 2.$
The volume of such $z$ is bounded by $N^{(n-1)(1-1/2n^{4})}$.
Otherwise, assume that $N^{-1/n}\beta(z)\geq N^{-1/2n}$. By minimality of $\beta(z)$, we have the bound
$\beta(z)\leq\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{z}=(dy_{1}\cdots y_{n-1})^{-1},$
using the computation of Remark 2.4. Therefore,
$dy_{1}\cdots y_{n-1}\leq N^{1/2n-1/n}=N^{-1/2n}.$
Since $y_{i}\gg 1$ for all $i\geq 2$, this also implies that
$dy_{1}\cdots y_{k}\ll N^{-1/2n}$
for all $1\leq k\leq n-1$.
Finally, by Lemma 2.7, we eliminate Case III, where we have shown that the minimal length in $L_{w}$ is $\ll N^{-1/n}$. Indeed, since Case III does not involve the Fricke involution, we preserve the assumption that $L_{w}$ is $\Omega$-balanced and arrive at a contradiction. ∎
Remark 5.3.
What we call the bulk $\Omega_{N}$ in the introduction of this paper is given here by case (2.b). The cuspidal regions that we remove are those corresponding to (1) and (2.a), as noted in (1.1).